Erin hires a maid to work in her home for$1,700 per month, the maid is 25 years old not related to Erin . during 2019 , the maid worked 5 months for Erin
1-what is the amount of social security tax Erin must pay as the maid's employer?
2- what is the amount of medicare tax Erin must pay as the maid's employer?
3- what is the amount Erin should withhold for social security and medicare tax?
Answer #1

Since wage is earned in 2019, taxes would be applicable for 2020.

Year 2020:

Social Security tax rate is 6.2% on the first $137,700 earned wages.

Medicare tax rate is1.45% on all the earned-wages.

Given,

Earned wages = $1,700 per month × 5 months

                        = 1,700 × 5

                        = $8,500

1.

Social Security tax (by employer) = Earned wages × 6.2%

                                                            = 8,500 × 6.2%

                                                            = $527 (Answer)

2.

Medicare tax (by employer) = Earned wages × 1.45%

                                                = 8,500 × 1.45%

                                                = $123.25 (Answer)

3.

Withholding amount = Earned wages × Total tax rates

                                    = 8,500 × (6.2% + 1.45%)

                                    = 8,500 × 7.65%

                                    = $650.25 (Answer)